OpenStudy (anonymous):

3 is what percent of 65

OpenStudy (lgbasallote):

just do \[\frac{3}{65} \times 100\]

OpenStudy (unklerhaukus):

three is what percent of 65 \[\Downarrow\]\[3~~~~~=w~\times~~~\frac{1}{100}~~\times 65\] \[\frac{3}{65}\times100=w\]